A rectangle with height and width equal to 3 and 25 respectively, is drawn on a checkered paper. Bazil paints a random horizontal 1×2 rectangle, and Peter paints a random vertical 2×1rectangle (each rectangle consists of 2 sells). Find the probability that at least one of the cells is painted twice. Express the answer in percent, and round to the nearest integer.
1
Expert's answer
2020-12-01T02:14:14-0500

"\text{Total number of horizontal rectangles 1*2:}"

"24*3 =72\text{ multiplying the number in a horizontal row by the number of rows}"

"\text{Total number of vertical rectangles 2*1:}"

"2*25=50 \text{ multiplying the number in a vertical row by the number of rows}"

"72*50=3600 \text{ of all possible combinations of the arrangement}"

"\text{of a pair of rectangles 1*2 and 2*1}"

"\text{Rectangles 1 * 2 and 2 * 1 intersect in a square 2 * 2 with 2 * 2 = 4 options}"

"\text{The number of squares 2 * 2 on an area of 3 * 25 will be:}"

"2*24 =48"

"\text{Number of intersections of rectangles1*2 and 2*1:}"

"48*4 =192"

"\text{Intersection probability percentage:}"

"\frac{192}{3600}*100\approx5"

Answer: 5% Intersection probability